OpenStudy (anonymous):

help please! 5X^3+25X^2-70X=0 solving by using both, Quadratic formula and Quadratic Equations. thank you!

OpenStudy (mathmate):

Hint: 5X^3+25X^2-70X=0 => \( 5x(x^2+5x−14)=0 \)

OpenStudy (anonymous):

thx!

OpenStudy (anonymous):

3X/4 -5/6=3/5 iam having a hard time finding the Common Dominator between 4, 6, and 5 ? help

OpenStudy (anonymous):

help! please.

OpenStudy (anonymous):

I GOT IT!!!

OpenStudy (anonymous):

how about ((X-3)/5))+((3/4))=((5/3))

OpenStudy (anonymous):

how about ((X-3)/5))+((3/4))=((5/3))

OpenStudy (mathmate):

Hint: multiply each term by 60 (LCM of 3,4,5) and work without fractions.